7. Given the joint density function /(x,y) =(kx (1 + 3 y*) 0<x<2,0<p?1 elsewhere a. Find k, g() h) and f(x) b. Evaluate P(-<X<1)
Answer #1

a)for above to be valid: \int_{0}^{2}\int_{0}^{1} f(x,y) dy dx =1

\int_{0}^{2}\int_{0}^{1} f(x,y) dy dx =\int_{0}^{2}\int_{0}^{1}kx(1+3y2) dy dx = \int_{0}^{2}kx(y+y3)|10 dx =\int_{0}^{2}kx(1+1) dx =2kx2/2|20 =4k =1

k=1/4

g(x) =\int_{0}^{1}f(x,y) dy=\int_{0}^{1}x(1+3y2)/4 dy =(x/4)*(y+y3)|10 =x/2

h(y) =\int_{0}^{2}f(x,y) dx=\int_{0}^{2}x(1+3y2)/4 dx =(x2/8)*(1+3y2)|20 =(1/2)*(1+3y2)

f(x|Y) =f(x,y)/h(y) =(1/4)x(1+3y2)/(1/2)*(1+3y2)) =x/2

b)P(1/4<X<1/2)=\int_{1/4}^{1/2} f(x) dx =\int_{1/4}^{1/2}x/2 dx = x2/4|1/21/4   =1/16-1/64 =3/64